# A free boundary problem in accretive growth

### v1

@@ 1.1. Accretive growth
Problem (1.1)–(1.2) is inspired by the modelization of an accretive-growth process in a confined environment. Accretive growth occurs across a variety of different biological systems, as well as in many natural and technological settings. The formation of horns, teeth and seashells [48, 54], coral reefs [35], bacterial colonies [30], trees [25], and cell motility due to actin growth [29] are biological examples of accreting systems, among many others [28]. In geophysics, sedimentation [12] and glacier evolution [32] are also accretive processes, so as planet formation [45]. Furthermore, accretive growth is a key aspect in metal solidification [46], crystal growth [38], additive manufacturing [31], layering, coating, and masonry, just to mention a few.
@@
Growth is always influenced by its environment and aspects such as nutrient or oxygen concentration, pressure, temperature, or stress, [28] have often to be taken into account. We model this fact by considering the simplest situation of a scalar activation field $u:Q_{v}\to{\mathbb{R}}$ driving the growth. The dynamics of the activation field $u$ is influenced by the actual shape of the growing system. Since growth is often very slow with respect to other processes, we assume that $u$ evolves quasistatically. We model a basic quasistatic diffusion process by the linear elliptic equation (1.2) on the growing domain $V(t)$ for all $t>0$. This corresponds to the case of $u$ being a concentration or temperature in an homogeneous medium and could of course be generalized. In particular, the choice of the right-hand side equal to $1$ is made solely for simplicity. Note that the resulting full model (1.1)–(1.2) is of free boundary type.
@@ 1.3. Results
The solution to the elliptic subproblem (1.2), (1.5)–(1.6) for a given $v$ relies on the possibility of establishing a Poincaré inequality on $V(t)$. In turn, this requires proving that $V(t)$ maintains some minimal level of regularity. In Proposition 3.8, we show that, starting from some suitably regular $V_{0}$, the growing set $V(t)$ is a John domain for all times (see Section 2). This extends to the constrained case $\Omega\not={\mathbb{R}}^{n}$ the arguments in [22]. Some technical modifications are required for this extension, since the boundary $\partial\Omega$ of the constraining domain must also be taken into account. Note that John regularity is sharp is this setting [22].
Eventually, the convergence of the iterative scheme depends on the time regularity of the set-valued maps $t\mapsto V(t)$ for different $v$. These happen to be Lipschitz continuous with respect to the Hausdorff distance, as well as Hölder continuous with respect to Lebesgue measure of the differences. Crucially, such continuities are uniform with respect to the iteration, allowing for the use of compactness arguments.
Let us conclude this section by noting that uniqueness of solutions to the coupled system (1.1)–(1.6) remains open. Even under the boundary condition (1.4), classical uniqueness arguments [5, 14, 40] seem to be not applicable. In fact, without further assumptions uniqueness seems to fails even in the simplest decoupled setting of $H$ not depending on $u$, see Remark 3.5 below.
Before moving on, let us mention that the results of this paper can be easily generalized to more complex settings. A first generalization consists in considering the Hamilton–Jacobi equation
$H(x,v(x),u(x,v(x)),\nabla v)=0\quad\in\text{in}\ \Omega\setminus\overline{V_{0}}$
@@ 1.4. Connections with previous studies
The theory of the Hamilton–Jacobi subproblem (1.1) with condition (1.3) for $u$ given (and suitably smooth) is well developed. We limit ourselves in mentioning the classical references [4, 5, 19, 40] for the general frame of viscosity solutions. In the context of optimal control with state constraints, condition (1.4) corresponds to the notion of constrained viscosity solution, going back to [14, 50, 51], see also [1, 26, 33, 36, 41] among many others. A class of biological invasion models based on Hamilton–Jacobi dynamics with obstacles has been introduced in [9] and studied in [10, 23], see also [7, 11] and the reference therein. Our results rely in a fundamental way on the metric structure underlying the Hamilton–Jacobi subproblem, which is already pointed out in [18, 40], see also [44, 47]. A wide range of results on the Hamilton–Jacobi subproblem is already available. However, to the best of our knowledge, none of them directly addresses the specific assumptions adopted here, see (3.5)–(3.9). For this reason, we provide detailed proofs.
@@
More generally, one can refer to the extensive literature on Hele–Shaw/Muskat/Stefan dynamics, see, e.g., [13, 55]. In these models, however, growth is driven by a fundamentally different underlying mechanism. In the Hele–Shaw setting, the driving field satisfies a Dirichlet boundary condition, and the domain evolves with a normal velocity proportional to the gradient of this field at the free boundary. In contrast, in our framework the activating field $u$ u is subject to a homogeneous Neumann condition, while the velocity of the free boundary depends on the Dirichlet values of $u$ at the growing front.
@@ 2.1. Assumptions
Given $x,\,y\in\overline{\Omega}$, we call geodesic from $y$ to $x$ any curve realizing the minimum above, recalling Note that geodesics are not necessarily unique. We additionally ask that the Euclidean and the geodesic distance are comparable in $\overline{\Omega}$, namely,
@@
$\Gamma\subset\partial V_{0}\cap\partial\Omega\ \text{nonempty, open in $\partial\Omega$, with $d_{H}(\Gamma,\partial V_{0}\setminus\partial\Omega)>0$}.$ (2.4)
As $V(t)\subset\Omega$ grows with $t$, assumption (2.4) guarantees that $\Gamma\subset\partial V(t)\cap\partial\Omega$ for all $t>0$ and $d_{H}(\Gamma,\partial V(t)\setminus\partial\Omega)$ is nondecreasing with respect to $t$.
As already mentioned, the open set $V(t)$ could be not smooth, see Proposition 3.8. Nonetheless, under assumptions (2.1) and (2.4), the portion $\Gamma$ of $\partial V(t)$ is smooth and is well separated from $\partial V(t)\setminus\partial\Omega$ for all $t>0$. This allows to define the trace on $\Gamma$ of any $u\in H^{1}(V(t))$. More precisely, set $G\subset\Gamma+B_{\delta}$ for some $\delta<d_{H}(\Gamma,\partial V_{0}\setminus\partial\Omega)$ small enough with $G\subset\subset V_{0}$ Lipschitz. Then, the trace operator on $H^{1}(V(t))$ can be defined by composing the restriction $u\in H^{1}(V(t))\mapsto u|_{G}\in H^{1}(G)$ and the trace operator $\gamma:H^{1}(G)\to H^{1/2}(\Gamma)$. This allows to specify
@@ 3. Some remarks on the Hamilton–Jacobi subproblem
Let us collect come basic properties of $\sigma$ in the following.
@@
$\quad\stackrel{{\scriptstyle p_{\zeta\delta}\in C_{x_{j}}}}{{\leq}}\liminf_{n\to\infty}\sigma(x_{j},q_{j})+\zeta+|q|\delta.$
@@
A consequence of Proposition 3.2 is that the minimum in (3.11) is always attained. In the following, we call optimal curve for $x$ a minimizer $\gamma_{x}\in{\rm Lip}_{x}$ in (3.11). Note that optimal curves $\gamma_{x}$ are not necessarily unique. Before moving on, let us mention that the assumptions of Proposition 3.2 could be somehow weakened (see [44]). We work nonetheless in the setting of (3.5)–(3.7) and (3.9), which allows the coupling with the PDE for the activating field $u$.
@@
$x\mapsto v_{0}(x)=(|x|-1)^{+}\quad\text{and}\quad x\mapsto v_{r}(x)=\min\{v_{0}(x),r-|x|\}\quad\forall r\geq 1$ (3.13)
@@
For all $x\in\overline{\Omega}$, $\overline{H}(x,\cdot)$ is the standard Minkowski functional of the convex set $C_{x}$, up to an additive constant. As $0\in B_{\sigma_{*}}\subset C_{x}$ we have that $\overline{H}(x,0)=-1$ and $p\mapsto\overline{H}(x,p)$ is positively $1$-homogeneous. Moreover, for all $(x,p)\in\overline{\Omega}\times{\mathbb{R}}^{n}$ with $H(x,p)\leq 0$ one has that $(x,p/1)\in C$, hence $\overline{H}(x,p)\leq 0$. On the other hand, if $\overline{H}(x,p)\leq 0$ there exists $\lambda_{\delta}\to 1$ such that $(x,p/\lambda_{\delta})\in C$ and the closure of $C$ from (3.5) ensures that $(x,p)\in C$, namely, $H(x,p)\leq 0$. This proves that $\{\overline{H}\leq 0\}=\{H\leq 0\}$. As assumption (3.8) directly implies that $\{H<0\}=\{\overline{H}<0\}$, by taking the complements we have that $\{\overline{H}\geq 0\}=\{H\geq 0\}$. We are hence allowed to replace $H$ by $\overline{H}$ in (3.2)–(3.3) without changing the corresponding viscosity solutions. Note more precisely that $C_{x}=\{p\in{\mathbb{R}}^{n}\ |\ \overline{H}(x,p)\leq 0\}$ for all $x\in\overline{\Omega}$, so that $\sigma$ and formula (3.14) remains unchanged.
@@
In case $\Omega$ is convex the proof of Proposition 3.3 can be simplified by skipping Step 1 and arguing directly on the extension $H_{\varepsilon}(x,p)=H(\pi(x),b_{\varepsilon}(x)p)$ where $\pi$ is the projection on $\overline{\Omega}$ and $b_{\varepsilon}(x)=a_{\varepsilon}(d(x,\overline{\Omega})/\varepsilon^{2})$ for some $a_{\varepsilon}:{\mathbb{R}}\to[\varepsilon,1]$ smooth with $a_{\varepsilon}(r)=1$ for $r\leq 1$ and $a_{\varepsilon}(r)=\varepsilon$ for $r\geq 2$. In fact, in this case one only needs $\pi$ to be a continuous map from ${\mathbb{R}}^{n}$ to $\overline{\Omega}$ coinciding with the identity on $\overline{\Omega}$. Hence, one could consider also some nonconvex $\overline{\Omega}=h(\overline{D})$ as indicated in Remark 3.4 by replacing $\pi$ in $H$ by $h\cdot\pi_{D}\circ h^{-1}$, where now $\pi_{D}$ is the projection on $D$.
@@
In case $\Omega={\mathbb{R}}^{n}$, the John regularity of $V(t)$ is proved in [22, Thm. 1.1]. In such unconstrained case, the assumptions on $H$ can be weakened and the constant $\kappa$ turns out to be depending just on $\kappa_{0}$ and, in particular, is independent of $T$. The constrained case $\Omega\not={\mathbb{R}}^{n}$ is more delicate, as the presence and the regularity of the boundary $\partial\Omega$ plays a role. Note nonetheless that John regularity is sharp for problem (3.2)–(3.3), see [22] for counterexamples.
@@
We divide the proof into steps. At first, we find an arc-length parametrized curve $\gamma$ in $\overline{\Omega}$ such that $d(\gamma(s),\partial V(t)\setminus\partial\Omega)\geq\overline{\kappa}s$ for some $\overline{\kappa}>0$ (Step 1, Figure 2). Then, starting from $\gamma$ we find a second curve $\widetilde{\gamma}$ such that $d(\widetilde{\gamma}(s),\partial V(t))\geq\widetilde{\kappa}s$ for some $0<\widetilde{\kappa}\leq\overline{\kappa}$ (Step 2, Figure 3). Eventually, we reparametrize $\widetilde{\gamma}$ by arc length and conclude that $V(t)$ is John domain (Step 3). The Lipschitz continuity of the set-valued map $t\mapsto V(t)$ with respect to the Hausdorff distance and the Hölder continuity with respect to the $L^{1}$ distance are proved in Step 4 and Step 5, respectively.
@@
see Figure 2. In particular, $\gamma(0)=x$, $\gamma(\ell_{\gamma})=x_{0}$, and $|\gamma^{\prime}|=1$ a.e. We devote the rest of this step to prove that, setting
@@
Moroever, one can prove that the ensuing $u:Q_{v}=\cup_{t>0}V(t)\times\{t\}\to{\mathbb{R}}$ is measurable. By choosing $w=u(\cdot,t)$ in (4.3) and using the uniform Poincaré inequality (4.2) one gets
@@
To proceed, we first first recall (3.32)–(3.33) from Proposition 3.8, which are uniform with respect to $v_{j}$, namely,

### v2

@@ 1.1. Accretive growth
Problem (1.1)–(1.2) is inspired by the modelization of an accretive-growth process in a confined environment. Accretive growth occurs across a variety of different biological systems, as well as in many natural and technological settings. The formation of horns, teeth, and seashells [48, 54], coral reefs [35], bacterial colonies [30], trees [25], and cell motility due to actin growth [29] are biological examples of accreting systems, among many others [28]. In geophysics, sedimentation [12] and glacier evolution [32] are also accretive processes, as well as planet formation [45]. Furthermore, accretive growth is a key aspect in metal solidification [46], crystal growth [38], additive manufacturing [31], layering, coating, and masonry, just to mention a few.
@@
Growth is always influenced by its environment and aspects such as nutrient or oxygen concentration, pressure, temperature, or stress, [28] have often to be taken into account. We model this fact by considering the simplest situation of a scalar activation field $u:Q_{v}\to{\mathbb{R}}$ driving the growth. The dynamics of the activation field $u$ is influenced by the actual shape of the growing system. Since growth is often very slow with respect to other processes, we assume that $u$ evolves quasistatically. We model a basic quasistatic diffusion process by the linear elliptic equation (1.2) on the growing domain $V(t)$ for all $t>0$. This corresponds to the case of $u$ being a concentration or temperature in a homogeneous medium and could of course be generalized. In particular, the choice of the right-hand side equal to $1$ is made solely for simplicity. Note that the resulting full model (1.1)–(1.2) is of free boundary type.
@@ 1.3. Results
The solution to the elliptic subproblem (1.2), (1.5)–(1.6) for a given $v$ relies on the possibility of establishing a Poincaré inequality on $V(t)$. In turn, this requires proving that $V(t)$ maintains some minimal level of regularity. In Proposition 3.8, we show that, starting from some suitably regular $V_{0}$, the growing set $V(t)$ is a John domain for all times (see Section 2). This extends to the constrained case $\Omega\not={\mathbb{R}}^{n}$ the arguments in [22]. Some technical modifications are required for this extension, since the boundary $\partial\Omega$ of the constraining domain must also be taken into account. Note that John regularity is sharp in this setting [22].
Eventually, the convergence of the iterative scheme depends on the time regularity of the set-valued maps $t\mapsto V(t)$ for different $v$. These are Lipschitz continuous with respect to the Hausdorff distance, as well as Hölder continuous with respect to Lebesgue measure of the differences. Crucially, such continuities are uniform with respect to the iteration, allowing for the use of compactness arguments.
Let us conclude this section by noting that uniqueness of solutions to the coupled system (1.1)–(1.6) remains open. Even under the boundary condition (1.4), classical uniqueness arguments [5, 14, 40] seem to be inapplicable. In fact, without further assumptions uniqueness seems to fail even in the simplest decoupled setting of $H$ not depending on $u$, see Remark 3.5 below.
Before moving on, let us mention that the results of this paper can be generalized to more complex settings. A first generalization consists in considering the Hamilton–Jacobi equation
$H(x,v(x),u(x,v(x)),\nabla v)=0\quad x\in\Omega\setminus\overline{V_{0}}$
@@ 1.4. Connections with previous studies
The theory of the Hamilton–Jacobi subproblem (1.1) with condition (1.3) for $u$ given (and suitably smooth) is well developed. We limit ourselves in mentioning the classical references [4, 5, 19, 40] for the general frame of viscosity solutions. In the context of optimal control with state constraints, condition (1.4) corresponds to the notion of constrained viscosity solution, going back to [14, 50, 51], see also [1, 26, 33, 36, 41] among many others. A class of biological invasion models based on Hamilton–Jacobi dynamics with obstacles has been introduced in [9] and studied in [10, 23], see also [7, 11] and the reference therein. Our results rely fundamentally on the metric structure underlying the Hamilton–Jacobi subproblem, which is already pointed out in [18, 40], see also [44, 47]. A wide range of results on the Hamilton–Jacobi subproblem is already available. However, to the best of our knowledge, none of them directly addresses the specific assumptions adopted here, see (3.5)–(3.9). For this reason, we provide detailed proofs.
@@
More generally, one can refer to the extensive literature on Hele–Shaw/Muskat/Stefan dynamics, see, e.g., [13, 55]. In these models, however, growth is driven by a fundamentally different underlying mechanism. In the Hele–Shaw setting, the driving field satisfies a Dirichlet boundary condition, and the domain evolves with a normal velocity proportional to the gradient of this field at the free boundary. In contrast, in our framework the activating field $u$ is subject to a homogeneous Neumann condition, while the velocity of the free boundary depends on the values of $u$ at the growing front.
@@ 2.1. Assumptions
Given $x,\,y\in\overline{\Omega}$, we call geodesic from $y$ to $x$ any curve realizing the minimum above, recalling that geodesics are not necessarily unique. We additionally ask that the Euclidean and the geodesic distance are comparable in $\overline{\Omega}$, namely,
@@
$\Gamma\subset\partial V_{0}\cap\partial\Omega\ \text{nonempty, open in $\partial\Omega$, with $\inf_{g\in\Gamma}d(g,\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\partial V_{0}\setminus\partial\Omega)>0$}.$ (2.4)
As $V(t)\subset\Omega$ grows with $t$, assumption (2.4) guarantees that $\Gamma\subset\partial V(t)\cap\partial\Omega$ for all $t>0$ and $\inf_{g\in\Gamma}d(g,\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\partial V_{0}\setminus\partial\Omega)>0$ is nondecreasing with respect to $t$.
As already mentioned, the open set $V(t)$ may be not smooth, see Proposition 3.8. Nonetheless, under assumptions (2.1) and (2.4), the portion $\Gamma$ of $\partial V(t)$ is smooth and is well separated from $\partial V(t)\setminus\partial\Omega$ for all $t>0$. This allows to define the trace on $\Gamma$ of any $u\in H^{1}(V(t))$. More precisely, set $G\subset\Gamma+B_{\delta}$ for some $\delta<\inf_{g\in\Gamma}d(g,\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\partial V_{0}\setminus\partial\Omega)>0$ small enough with $G\subset\subset V_{0}$ Lipschitz. Then, the trace operator on $H^{1}(V(t))$ can be defined by composing the restriction $u\in H^{1}(V(t))\mapsto u|_{G}\in H^{1}(G)$ and the trace operator $\gamma:H^{1}(G)\to H^{1/2}(\Gamma)$. This allows to specify
@@ 3. Some remarks on the Hamilton–Jacobi subproblem
Let us collect some basic properties of $\sigma$ in the following.
@@
$\quad\stackrel{{\scriptstyle p_{\zeta\delta}\in C_{x_{j}}}}{{\leq}}\liminf_{j\to\infty}\sigma(x_{j},q_{j})+\zeta+|q|\delta.$
@@
A consequence of Proposition 3.2 is that the minimum in (3.11) is always attained. In the following, we call optimal curve for $x$ a minimizer $\gamma_{x}\in{\rm Lip}_{x}$ in (3.11). Note that optimal curves $\gamma_{x}$ are not necessarily unique. Before moving on, let us mention that the assumptions of Proposition 3.2 could be relaxed (see [44]). We work nonetheless in the setting of (3.5)–(3.7) and (3.9), which allows the coupling with the PDE for the activating field $u$.
@@
$x\mapsto v_{0}(x)=(|x|-1)^{+}\quad\text{and}\quad x\mapsto v_{r}(x)=\min\{v_{0}(x),r-|x|\}\quad\forall r\geq 1,$ (3.13)
@@
For all $x\in\overline{\Omega}$, $\overline{H}(x,\cdot)$ is the standard Minkowski functional of the convex set $C_{x}$, up to an additive constant. As $0\in B_{\sigma_{*}}\subset C_{x}$ we have that $\overline{H}(x,0)=-1$ and $p\mapsto\overline{H}(x,p)$ is positively $1$-homogeneous. Moreover, for all $(x,p)\in\overline{\Omega}\times{\mathbb{R}}^{n}$ with $H(x,p)\leq 0$ one has that $(x,p/1)\in C$, hence $\overline{H}(x,p)\leq 0$. On the other hand, if $\overline{H}(x,p)\leq 0$ there exists $\lambda_{\delta}\to 1$ such that $(x,p/\lambda_{\delta})\in C$ and the closure of $C$ from (3.5) ensures that $(x,p)\in C$, namely, $H(x,p)\leq 0$. This proves that $\{\overline{H}\leq 0\}=\{H\leq 0\}$. As assumption (3.8) directly implies that $\{H<0\}=\{\overline{H}<0\}$, by taking the complements we have that $\{\overline{H}\geq 0\}=\{H\geq 0\}$. We are hence allowed to replace $H$ by $\overline{H}$ in (3.2)–(3.3) without changing the corresponding viscosity solutions. Note more precisely that $C_{x}=\{p\in{\mathbb{R}}^{n}\ |\ \overline{H}(x,p)\leq 0\}$ for all $x\in\overline{\Omega}$, so that $\sigma$ and formula (3.14) remain unchanged.
@@
In case $\Omega$ is convex the proof of Proposition 3.3 can be simplified by skipping Step 1 and arguing directly on the extension $H_{\varepsilon}(x,p)=H(\pi(x),b_{\varepsilon}(x)p)$ where $\pi$ is the projection on $\overline{\Omega}$ and $b_{\varepsilon}(x)=a_{\varepsilon}(d(x,\overline{\Omega})/\varepsilon^{2})$ for some $a_{\varepsilon}:{\mathbb{R}}\to[\varepsilon,1]$ smooth with $a_{\varepsilon}(r)=1$ for $r\leq 1$ and $a_{\varepsilon}(r)=\varepsilon$ for $r\geq 2$. In fact, in this case one only needs $\pi$ to be a continuous map from ${\mathbb{R}}^{n}$ to $\overline{\Omega}$ coinciding with the identity on $\overline{\Omega}$. Hence, one could also consider some nonconvex $\overline{\Omega}=h(\overline{D})$ as indicated in Remark 3.4 by replacing $\pi$ in $H$ by $h\circ\pi_{D}\circ h^{-1}$, where now $\pi_{D}$ is the projection on $D$.
@@
In case $\Omega={\mathbb{R}}^{n}$, the John regularity of $V(t)$ is proved in [22, Thm. 1.1]. In such unconstrained case, the assumptions on $H$ can be weakened and the constant $\kappa$ turns out to be depending only on $\kappa_{0}$ and, in particular, is independent of $T$. The constrained case $\Omega\not={\mathbb{R}}^{n}$ is more delicate, as the presence and the regularity of the boundary $\partial\Omega$ plays a role. Note nonetheless that John regularity is sharp for problem (3.2)–(3.3), see [22] for counterexamples.
@@
We divide the proof into steps. At first, we find an arc-length parametrized curve $\gamma$ in $\overline{\Omega}$ such that $d(\gamma(s),\partial V(t)\setminus\partial\Omega)\geq\overline{\kappa}s$ for some $\overline{\kappa}>0$ (Step 1, Figure 2). Then, starting from $\gamma$ we find a second curve $\widetilde{\gamma}$ such that $d(\widetilde{\gamma}(s),\partial V(t))\geq\widetilde{\kappa}s$ for some $0<\widetilde{\kappa}\leq\overline{\kappa}$ (Step 2, Figure 3). Eventually, we reparametrize $\widetilde{\gamma}$ by arc length and conclude that $V(t)$ is a John domain (Step 3). The Lipschitz continuity of the set-valued map $t\mapsto V(t)$ with respect to the Hausdorff distance and the Hölder continuity with respect to the $L^{1}$ distance are proved in Step 4 and Step 5, respectively.
@@
see Figure 2. In particular, $\gamma(0)=x$, $\gamma(\ell_{\gamma})=x_{0}$, and $|\gamma^{\prime}|=1$ a.e. We devote the rest of this step to proving that, setting
@@
Moreover, one can prove that the ensuing $u:Q_{v}=\cup_{t>0}V(t)\times\{t\}\to{\mathbb{R}}$ is measurable. By choosing $w=u(\cdot,t)$ in (4.3) and using the uniform Poincaré inequality (4.2) one gets
@@
To proceed, we first recall (3.32)–(3.33) from Proposition 3.8, which are uniform with respect to $v_{j}$, namely,